OpenStudy (anonymous):

solve the compound inequality -16 ≤ 2 + 9x ≤ 11

OpenStudy (ranga):

Isolate the x first: Subtract 2 from left, center , right \[-18 \le 9x \le 9\] Divide throughout by 9 \[-2 \le x \le 1\]